OpenStudy (anonymous):

line AB represented by the equation y=3x-2 line CD is parallel to line AB and passes through thepoint 5,2 what is the equation of line CD in slope intercept form?

OpenStudy (anonymous):

because it is parallel, it has the same slope: 3 so using point slope form, y-2=3(x-5) which is y=3x-13